Mein Name

Meine Straße 1 12345 Meindorf

Mein Name · Meine Straße 1 · D–12345 Meindorf

Einschreiben

DANTE e.V. Postfach 101840 69120 Heidelberg

A. Kielhorn

Telephone: 89 34 39

e-mail: a.kielhorn@tu-bs.de

1.6.1994^{Your letter of :}  dante^{Your ref. :} Subject: LaTeX2e Letter-style  March 3, 1995

Dear friends,

This is a demonstration-letter for my new myletter-class. It is based on letter.cls and liletter (Stefan Lindner Letter Style, distributed with Atari TEX).

It supports everything needed like your-ref, your-mail subject and of course configurable name and signature.

Everything site-depending like header, footer or returnaddress went into a .cfg-file. The \usename command selects a configuration file and enables you do have different information for business and privat letters. You may even set up a .cfg-file for your husband.

Keep TEXing

(Axel Kielhorn)

cc : de.comp.tex comp.text.tex

encl. : myletter.cls default.cfg

P.S. see default.cfg for more information